OpenStudy (crashonce):

More @jim_thompson5910 sorry

OpenStudy (crashonce):

\[\frac{ 1+3+5+...+(2n-1) }{ 2+4+6+...+2n} = \frac{ 115 }{ 116 }\]

OpenStudy (crashonce):

find a positive integral solution for n

OpenStudy (crashonce):

choices: 110, 115, 116, 231, or no answer

jimthompson5910 (jim_thompson5910):

again I'm going to use 1+2+3+...+n = n*(n+1)/2

jimthompson5910 (jim_thompson5910):

1+3+5+...+(2n-1) = 2*1-1+2*2-1+2*3-1+...+2*n-1 1+3+5+...+(2n-1) = (2*1+2*2+2*3+...+2*n) + (-1 - 1 - 1...-1) 1+3+5+...+(2n-1) = 2*(1+2+3+...+n) + (-1 - 1 - 1...-1) 1+3+5+...+(2n-1) = 2*n*(n+1)/2 + n*(-1) 1+3+5+...+(2n-1) = n*(n+1) - n 2+4+6+...+2n = 2*(1+2+3+...+n) 2+4+6+...+2n = 2*n*(n+1)/2 2+4+6+...+2n = n*(n+1)

jimthompson5910 (jim_thompson5910):

the numerator simplifies to n*(n+1) - n the denominator simplifies to n*(n+1)

jimthompson5910 (jim_thompson5910):

and I guess you could factor out n to get n*(n+1) - n = n[(n+1) - 1] n*(n+1) - n = n[n+1 - 1] n*(n+1) - n = n[n+0] n*(n+1) - n = n*n n*(n+1) - n = n^2

OpenStudy (anonymous):

It is an arithmetic progression divided by another arithmetic progression.

jimthompson5910 (jim_thompson5910):

\[\large \frac{ 1+3+5+...+(2n-1) }{ 2+4+6+...+2n} = \frac{ 115 }{ 116 }\] turns into \[\large \frac{ n^2 }{ n*(n+1)} = \frac{ 115 }{ 116 }\] which simplifies to \[\large \frac{ n }{ n+1} = \frac{ 115 }{ 116 }\]

OpenStudy (anonymous):

So you could use: \[ \sum a_n=\frac n2 (a_i+a_n) \]

jimthompson5910 (jim_thompson5910):

yes that's probably a much faster way

OpenStudy (crashonce):

so its 115

jimthompson5910 (jim_thompson5910):

yeah